# Rep3D: Re-parameterize Large 3D Kernels with Low-Rank Receptive Modeling for Medical Imaging

### v1

@@ 3.2 Low-Rank Receptive Bias Modeling (LRBM)
As the above theory further validates the correlation between variable learning with the local-to-global gradient dynamics in ERF, we argue that such receptive bias can enhance the efficiency of learning large convolution kernels. We model the diffusion behavior of ERF with a reciprocal distance decay function $f_{d}$ and generate a prior mapping $P\in\mathbb{R}^{C\times 1\times K\times K\times K}$ for weight re-parameterization as follow:
$f_{d}(x,y,z,c)$ $=\sqrt{\left({x-c}\right)^{2}+\left({y-c}\right)^{2}+\left({z-c}\right)^{2}}$ (14)
$P$ $=\frac{\beta}{d(x_{k},y_{k},z_{k},c)+\beta}$
where $k$ and $c$ are the element and central index of the kernel weight, $\beta$ is a learnable parameter to control the weight distribution of the distance mapping and initialize as 0. However, such a fixed prior mapping lacks of flexibility to adapt the weighting importance dynamically across the fine-grained semantic variations in medical imaging. To address this, we propose to adapt learnable spatial bias by co-training a light-weight 2-layer generator network $f_{\theta}:\mathbb{R}^{C\times 1\times K\times K\times K}\rightarrow\mathbb{R}^{C\times 1\times K\times K\times K}$. We generate an adaptive mask $M$ for depthwise convolution kernels with low computation cost as follows:
$M=P+f_{\theta}(P)$ (15)
$f_{\theta}(P)=\text{Norm}_{2}(\text{DConv}_{2}\left(\sigma\left(\text{Norm}_{1}(\text{DConv}_{1}(P))\right)\right))$ (16)
where $\text{DConv}_{1}$ and $\text{DConv}_{2}$ are 3D depthwise convolutions with kernel size of 7 and padding of 3, both $\text{Norm}_{1}$ $\text{Norm}_{2}$ are the layer normalizations, and $\sigma$ is a non-linear sigmoid activation to ensure all scaling value between 0 and 1. Such learnable function aims to capture the dynamic weighting of each kernel elements across local to global, while preserving computational efficiency. The resulting modulation mask $M$ is then used to reparameterize the kernel weights:
$W_{\text{eff}}=W\odot M$ (17)
where $W$ is the original convolution kernel and $\odot$ denotes element-wise multiplication. Importantly, the mask is applied during training only and the learned generator can be removed during inference for efficiency.
@@ 3.3 Network Architecture
The overall network architecture to validate Rep3D builds upon the encoder–decoder structure of 3D UX-Net [18], which processes volumetric data through hierarchical resolution stages with skip connections to preserve fine-grained spatial features. Unlike prior transformer-based models or heavily modular CNNs, our design favors plain convolution blocks to minimize computational burden while preserving capacity for large-scale context modeling. Following the insights from prior work [7], we adopt a $21\times 21\times 21$ depthwise convolution (DWC-21) as the kernel backbone, which we empirically identify as the best trade-off between expressiveness and efficiency in 3D. Each encoder block consists of batch normalization, followed by the depthwise convolution and GeLU activation. The feature propagation from layer $\ell-1$ to $\ell$ and then to $\ell+1$ is defined as:
$\hat{z}_{\ell}=\text{GELU}\left(\text{DWC-21}(\text{BN}(z_{\ell-1}))\right),\quad\hat{z}_{\ell+1}=\text{GELU}\left(\text{DWC-21}(\text{BN}(\hat{z}_{\ell}))\right)$ (18)
where $z_{\ell-1}$ is the input from the previous layer, $\hat{z}\ell$ and $\hat{z}{\ell+1}$ are intermediate representations, BN denotes batch normalization, and DWC-21 represents depthwise convolution with a $21^{3}$ kernel. This architectural choice allows the network to efficiently encode both local and global context, while enabling seamless integration of our re-parameterized learning framework (as detailed in section 3.1 and 3.2). The simplicity of the block ensures compatibility with the spatial modulation mask described in the next section and avoids unnecessary overhead during both training and inference.
@@ 4 Experimental Setup
Datasets. We evaluate Rep3D on four publicly available volumetric segmentation datasets, covering a wide range of anatomical structures across different spatial scales—from large organs (e.g., liver, stomach) to smaller and more challenging targets (e.g., tumors, vessels): 1) AMOS22 (MICCAI 2022 Abdominal Multi-organ Segmentation Challenge) [15]: Comprises 200 multi-contrast abdominal CT scans with 15 organ-level anatomical labels and 33 MRI scans with 13 organ-level anatomical labels for comprehensive abdominal segmentation, 2) KiTS19 (MICCAI 2019 Kidney Tumor Segmentation Challenge) [12]: Includes 210 contrast-enhanced abdominal CT scans from the University of Minnesota Medical Center (2010–2018), with manual annotations for kidney, tumor, and cyst, 3) MSD Pancreas (Medical Segmentation Decathlon) [1]: Contains 282 abdominal contrast-enhanced CT scans annotated for both pancreas and pancreatic tumor segmentation, and 4) MSD Hepatic Vessel (Medical Segmentation Decathlon) [1]: Contains 303 abdominal CT scans annotated for hepatic vessel and associated tumor segmentation. Additional dataset details, including resolution normalization, voxel spacing, and pre-processing pipelines, are provided in the supplementary material.
Implementation Details. All experiments are conducted under a direct supervised learning setting. For the KiTS and MSD datasets, we employ a 5-fold cross-validation strategy using an 80%/10%/10% split for training, validation, and testing, respectively. For the AMOS dataset, we use a fixed single split with the same partitioning ratio. Details on training procedures and preprocessing protocols are provided in the supplementary material. Our proposed re-parameterization approach Rep3D, is benchmarked against both convolutional and transformer-based state-of-the-art (SOTA) methods for 3D medical image segmentation. For nnUNet [14], we evaluate performance across two training schedules to account for fairness, since Rep3D is trained with 60,000 iterations (approximately equivalent to 350 epochs). We report results using the Dice Similarity Coefficient (DSC) as the primary evaluation metric, quantifying spatial overlap between predicted segmentations and ground truth labels. Additionally, we conduct ablation studies to analyze: (1) the effect of kernel size in the spatial bias generator, and (2) the adaptability of Rep3D when integrated into other 3D large kernel architectures (e.g., 3D UX-Net).
@@
KiTS MSD Methods #Params FLOPs Kidney Tumor Cyst Mean Pancreas Tumor Mean Hepatic Tumor Mean 3D U-Net [5] 4.81M 135.9G 0.918 0.657 0.361 0.645 0.711 0.584 0.648 0.569 0.609 0.589 SegResNet [26] 1.18M 15.6G 0.935 0.713 0.401 0.683 0.740 0.613 0.677 0.620 0.656 0.638 RAP-Net [20] 38.2M 101.2G 0.931 0.710 0.427 0.689 0.742 0.621 0.682 0.610 0.643 0.627 nn-UNet [14] 31.2M 743.3G 0.943 0.732 0.443 0.706 0.775 0.630 0.703 0.623 0.695 0.660 TransBTS [28] 31.6M 110.3G 0.932 0.691 0.384 0.669 0.749 0.610 0.679 0.589 0.636 0.613 UNETR [11] 92.8M 82.5G 0.921 0.669 0.354 0.648 0.735 0.598 0.667 0.567 0.612 0.590 nnFormer [32] 149.3M 213.0G 0.930 0.687 0.376 0.664 0.769 0.603 0.686 0.591 0.635 0.613 SwinUNETR [10] 62.2M 328.1G 0.939 0.702 0.400 0.680 0.785 0.632 0.708 0.622 0.647 0.635 3D UX-Net (k=7) [18] 53.0M 639.4G 0.942 0.724 0.425 0.697 0.737 0.614 0.676 0.625 0.678 0.652 UNesT-B [31] 87.2M 258.4G 0.943 0.746 0.451 0.710 0.778 0.601 0.690 0.611 0.645 0.640 Rep3D (Fixed Prior) 65.8M 757.4G 0.950 0.757 0.473 0.727 0.789 0.640 0.715 0.635 0.681 0.658 Rep3D 66.0M 757.6G 0.955 0.763 0.490 0.736* 0.793 0.653 0.723* 0.650 0.697 0.674*
Table 2: Evaluations on the AMOS testing split in different scenarios.(*: $p<0.01$, with Paired Wilcoxon signed-rank test to all baseline networks)
@@
AMOS CT (Train From Scratch Scenario) Methods Spleen R. Kid L. Kid Gall. Eso. Liver Stom. Aorta IVC Panc. RAG LAG Duo. Blad. Pros. Avg nn-UNet (350 Epochs) 0.951 0.919 0.930 0.845 0.797 0.975 0.863 0.941 0.898 0.813 0.730 0.677 0.772 0.797 0.815 0.850 nn-UNet (1000 Epochs) 0.967 0.958 0.945 0.890 0.818 0.979 0.914 0.953 0.920 0.824 0.799 0.743 0.823 0.900 0.867 0.887 TransBTS 0.930 0.921 0.909 0.798 0.722 0.966 0.801 0.900 0.820 0.702 0.641 0.550 0.684 0.730 0.679 0.783 UNETR 0.925 0.923 0.903 0.777 0.701 0.964 0.759 0.887 0.821 0.687 0.688 0.543 0.629 0.710 0.707 0.740 nnFormer 0.932 0.928 0.914 0.831 0.743 0.968 0.820 0.905 0.838 0.725 0.678 0.578 0.677 0.737 0.596 0.785 SwinUNETR 0.956 0.957 0.949 0.891 0.820 0.978 0.880 0.939 0.894 0.818 0.800 0.730 0.803 0.849 0.819 0.871 3D UX-Net (k=7) 0.966 0.959 0.951 0.903 0.833 0.980 0.910 0.950 0.913 0.830 0.805 0.756 0.846 0.897 0.863 0.890 3D UX-Net (k=21) 0.963 0.959 0.953 0.921 0.848 0.981 0.903 0.953 0.910 0.828 0.815 0.754 0.824 0.900 0.878 0.891 UNesT-B 0.966 0.961 0.956 0.903 0.840 0.980 0.914 0.947 0.912 0.838 0.803 0.758 0.846 0.895 0.854 0.891 RepOptimizer 0.968 0.964 0.953 0.903 0.857 0.981 0.915 0.950 0.915 0.826 0.802 0.756 0.813 0.906 0.867 0.892 Rep3D (Fixed Prior) 0.972 0.963 0.964 0.911 0.861 0.982 0.921 0.956 0.924 0.837 0.818 0.777 0.831 0.916 0.879 0.902 Rep3D (LRBM) 0.978 0.970 0.964 0.928 0.871 0.984 0.927 0.960 0.930 0.851 0.828 0.784 0.850 0.920 0.881 0.910* AMOS MRI (Train From Scratch Scenario) Methods Spleen R. Kid L. Kid Gall. Eso. Liver Stom. Aorta IVC Panc. RAG LAG Duo. Blad. Pros. Avg nn-UNet (350 Epochs) 0.967 0.855 0.958 0.663 0.736 0.973 0.888 0.956 0.907 0.793 0.533 0.572 0.668 - - 0.805 nn-UNet (1000 Epochs) 0.973 0.940 0.965 0.681 0.810 0.980 0.893 0.967 0.917 0.834 0.667 0.689 0.701 - - 0.847 TransBTS 0.956 0.957 0.955 0.619 0.770 0.974 0.867 0.958 0.852 0.836 0.591 0.630 0.648 - - 0.816 UNETR 0.942 0.956 0.930 0.552 0.741 0.967 0.836 0.947 0.829 0.815 0.564 0.621 0.624 - - 0.794 nnFormer 0.949 0.952 0.950 0.601 0.758 0.972 0.859 0.960 0.843 0.832 0.569 0.618 0.637 - - 0.808 SwinUNETR 0.972 0.961 0.961 0.649 0.814 0.978 0.889 0.961 0.862 0.854 0.659 0.649 0.664 - - 0.836 3D UX-Net (k=7) 0.971 0.965 0.966 0.603 0.828 0.978 0.869 0.962 0.878 0.837 0.696 0.689 0.696 - - 0.841 3D UX-Net (k=21) 0.968 0.962 0.967 0.610 0.830 0.977 0.858 0.954 0.880 0.829 0.701 0.697 0.700 - - 0.840 UNesT-B 0.971 0.965 0.967 0.615 0.831 0.980 0.865 0.949 0.883 0.845 0.852 0.700 0.697 - - 0.854 RepOptimizer 0.970 0.967 0.971 0.635 0.823 0.978 0.875 0.963 0.882 0.850 0.689 0.691 0.711 - - 0.847 Rep3D (Fixed Prior) 0.972 0.965 0.970 0.644 0.838 0.980 0.883 0.965 0.893 0.861 0.714 0.701 0.725 - - 0.855 Rep3D (LRBM) 0.975 0.969 0.975 0.657 0.845 0.984 0.891 0.970 0.901 0.879 0.718 0.721 0.750 - - 0.864*
@@ 5.1 Evaluation on Tissue & Tumor Segmentation
To assess the generalization and scalability of Rep3D across diverse anatomical structures and clinical targets, we evaluate performance on three representative volumetric segmentation tasks using the KiTS, MSD Pancreas, and MSD Hepatic Vessel datasets. As shown in Table 3, Rep3D achieves state-of-the-art performance across all settings, consistently outperforming both convolution- and transformer-based baselines. On the KiTS dataset, which includes kidney, tumor, and cyst segmentation, Rep3D achieves the highest average Dice score of 0.736, with strong individual scores of 0.955 (kidney), 0.763 (tumor), and 0.490 (cyst). Notably, Rep3D improves tumor segmentation performance by 2.28% Dice over UNesT-B and 5.39% Dice over 3D UX-Net, demonstrating its ability to adapt to complex local variations in pathological regions. On the MSD Pancreas task, which is particularly challenging due to the pancreas’s low contrast and irregular boundaries, Rep3D sets a new benchmark with an average Dice score of 0.723, outperforming SwinUNETR (0.708), nnUNet (0.703), and UNesT-B (0.690). Tumor segmentation also benefits from our re-parameterization design, improving by 3.32% Dice compared to 3D UX-Net and 2,03% Dice compared to the fixed-prior variant. On the MSD Hepatic Vessel dataset, Rep3D continues to lead with a mean Dice of 0.674, outperforming the previous best model (UNesT-B, 0.640) and demonstrating superior vessel and tumor localization. The results also highlight the effectiveness of Rep3D’s spatially adaptive learning dynamics, especially in sparse and small-structure segmentation where traditional large-kernel convolutions or global self-attention tend to underperform.

### v2

@@ 3.2 Low-Rank Receptive Bias Modeling (LRBM)
As the above theory validates the correlation between variable learning and local-to-global gradient dynamics in ERF, we argue that such receptive bias can enhance the efficiency of learning large convolution kernels. We model the diffusion behavior of ERF with a reciprocal distance decay function $f_{d}$ and generate a prior mapping $P\in\mathbb{R}^{C\times 1\times K\times K\times K}$ for weight re-parameterization as follows:
$f_{d}(x,y,z,c)$ $=\sqrt{(x-c)^{2}+(y-c)^{2}+(z-c)^{2}},$ (12)
$P$ $=\frac{\beta}{f_{d}(x_{k},y_{k},z_{k},c)+\beta},$ (13)
where $k$ and $c$ are the element and central index of the kernel weight, and $\beta$ is a learnable parameter initialized to 0 that controls the weight distribution of the distance mapping. However, such a fixed prior mapping lacks the flexibility to adapt the weighting importance dynamically across the fine-grained semantic variations in medical imaging. To address this, we propose to adapt learnable spatial bias by co-training a lightweight 2-layer generator network $f_{\theta}:\mathbb{R}^{C\times 1\times K\times K\times K}\rightarrow\mathbb{R}^{C\times 1\times K\times K\times K}$. We generate an adaptive mask $M$ for depthwise convolution kernels with low computation cost as follows:
$M=P+f_{\theta}(P),$ (14)
with the generator defined as:
$f_{\theta}(P)=\mathrm{Norm}_{2}\Big(\mathrm{DConv}_{2}\big(\sigma(\mathrm{Norm}_{1}(\mathrm{DConv}_{1}(P)))\big)\Big),$ (15)
where $\mathrm{DConv}_{1}$ and $\mathrm{DConv}_{2}$ are 3D depthwise convolutions with kernel size 7 and padding 3, $\mathrm{Norm}_{1}$ and $\mathrm{Norm}_{2}$ are layer normalizations, and $\sigma$ is a sigmoid activation ensuring all scaling values are between 0 and 1. Such a learnable function aims to capture the dynamic weighting of each kernel element across local to global contexts while preserving computational efficiency. The resulting modulation mask $M$ is then used to re-parameterize the kernel weights:
$W_{\mathrm{eff}}=W\odot M,$ (16)
where $W$ is the original convolution kernel and $\odot$ denotes element-wise multiplication. Importantly, the mask is applied during training only, and the learned generator can be removed during inference for efficiency.
@@ 3.3 Network Architecture
The overall network architecture to validate Rep3D builds upon the encoder–decoder structure of 3D UX-Net (Lee et al., 2022), which processes volumetric data through hierarchical resolution stages with skip connections to preserve fine-grained spatial features. Unlike prior transformer-based models or heavily modular CNNs, our design favors plain convolution blocks to minimize computational burden while preserving capacity for large-scale context modeling. Following insights from prior work (Ding et al., 2022b), we adopt a $21\times 21\times 21$ depthwise convolution (DWC-21) as the kernel backbone, which we empirically identify as the best trade-off between expressiveness and efficiency in 3D. Each encoder block consists of batch normalization, followed by the depthwise convolution and GELU activation. The feature propagation from layer $\ell-1$ to layer $\ell$ and then to $\ell+1$ is defined as:
$\hat{z}_{\ell}$ $=\mathrm{GELU}\big(\mathrm{DWC}_{21}(\mathrm{BN}(z_{\ell-1}))\big),$ (17)
[…]
where $z_{\ell-1}$ is the input from the previous layer, $\hat{z}_{\ell}$ and $\hat{z}_{\ell+1}$ are intermediate representations, $\mathrm{BN}$ denotes batch normalization, and $\mathrm{DWC}_{21}$ represents depthwise convolution with a $21^{3}$ kernel. This architectural choice allows the network to efficiently encode both local and global context, while enabling seamless integration of our re-parameterized learning framework.
@@ 4 Experimental Setup
Datasets and Implementation Details. We evaluate Rep3D on five publicly available volumetric segmentation datasets, covering a wide range of anatomical structures across different spatial scales, from large organs (e.g., liver, stomach) to smaller and more challenging targets (e.g., tumors, vessels). We report results using the Dice Similarity Coefficient (DSC) as the primary evaluation metric, quantifying spatial overlap between predicted segmentations and ground truth labels. Additional details, including dataset resolution normalization, voxel spacing, and pre-processing pipelines and experimental details are provided in the appendix.
@@
KiTS MSD Pancreas MSD Hepatic Methods #Params FLOPs Kidney Tumor Cyst Mean Pancreas Tumor Mean Hepatic Tumor Mean 3D U-Net (Çiçek et al., 2016) 4.81M 135.9G 0.918 0.657 0.361 0.645 0.711 0.584 0.648 0.569 0.609 0.589 SegResNet (Myronenko, 2018) 1.18M 15.6G 0.935 0.713 0.401 0.683 0.740 0.613 0.677 0.620 0.656 0.638 RAP-Net (Lee et al., 2021) 38.2M 101.2G 0.931 0.710 0.427 0.689 0.742 0.621 0.682 0.610 0.643 0.627 nn-UNet (Isensee et al., 2021) 31.2M 743.3G 0.943 0.732 0.443 0.706 0.775 0.630 0.703 0.623 0.695 0.660 TransBTS (Wang et al., 2021) 31.6M 110.3G 0.932 0.691 0.384 0.669 0.749 0.610 0.679 0.589 0.636 0.613 UNETR (Hatamizadeh et al., 2022b) 92.8M 82.5G 0.921 0.669 0.354 0.648 0.735 0.598 0.667 0.567 0.612 0.590 nnFormer (Zhou et al., 2021) 149.3M 213.0G 0.930 0.687 0.376 0.664 0.769 0.603 0.686 0.591 0.635 0.613 SwinUNETR (Hatamizadeh et al., 2022a) 62.2M 328.1G 0.939 0.702 0.400 0.680 0.785 0.632 0.708 0.622 0.647 0.635 3D UX-Net (k=7) (Lee et al., 2022) 53.0M 639.4G 0.942 0.724 0.425 0.697 0.737 0.614 0.676 0.625 0.678 0.652 UNesT-B (Yu et al., 2023) 87.2M 258.4G 0.943 0.746 0.451 0.710 0.778 0.601 0.690 0.611 0.645 0.640 Rep3D (Fixed Prior) 65.8M 757.4G 0.950 0.757 0.473 0.727 0.789 0.640 0.715 0.635 0.681 0.658 Rep3D 66.0M 757.6G 0.955 0.763 0.490 0.736* 0.793 0.653 0.723* 0.650 0.697 0.674*
Table 2: Evaluations on the AMOS testing split in different scenarios (*: $p<0.01$).
@@
AMOS CT (Train From Scratch Scenario) Methods Spleen R.Kid L.Kid Gall. Eso. Liver Stom. Aorta IVC Panc. RAG LAG Duo. Blad. Pros. Avg nn-UNet (350 Ep) 0.951 0.919 0.930 0.845 0.797 0.975 0.863 0.941 0.898 0.813 0.730 0.677 0.772 0.797 0.815 0.850 nn-UNet (1000 Ep) 0.967 0.958 0.945 0.890 0.818 0.979 0.914 0.953 0.920 0.824 0.799 0.743 0.823 0.900 0.867 0.887 TransBTS 0.930 0.921 0.909 0.798 0.722 0.966 0.801 0.900 0.820 0.702 0.641 0.550 0.684 0.730 0.679 0.783 UNETR 0.925 0.923 0.903 0.777 0.701 0.964 0.759 0.887 0.821 0.687 0.688 0.543 0.629 0.710 0.707 0.740 nnFormer 0.932 0.928 0.914 0.831 0.743 0.968 0.820 0.905 0.838 0.725 0.678 0.578 0.677 0.737 0.596 0.785 SwinUNETR 0.956 0.957 0.949 0.891 0.820 0.978 0.880 0.939 0.894 0.818 0.800 0.730 0.803 0.849 0.819 0.871 3D UX-Net (k=7) 0.966 0.959 0.951 0.903 0.833 0.980 0.910 0.950 0.913 0.830 0.805 0.756 0.846 0.897 0.863 0.890 3D UX-Net (k=21) 0.963 0.959 0.953 0.921 0.848 0.981 0.903 0.953 0.910 0.828 0.815 0.754 0.824 0.900 0.878 0.891 UNesT-B 0.966 0.961 0.956 0.903 0.840 0.980 0.914 0.947 0.912 0.838 0.803 0.758 0.846 0.895 0.854 0.891 RepOptimizer 0.968 0.964 0.953 0.903 0.857 0.981 0.915 0.950 0.915 0.826 0.802 0.756 0.813 0.906 0.867 0.892 Rep3D (Fixed) 0.972 0.963 0.964 0.911 0.861 0.982 0.921 0.956 0.924 0.837 0.818 0.777 0.831 0.916 0.879 0.902 Rep3D (LRBM) 0.978 0.970 0.964 0.928 0.871 0.984 0.927 0.960 0.930 0.851 0.828 0.784 0.850 0.920 0.881 0.910* AMOS MRI (Train From Scratch Scenario) Methods Spleen R.Kid L.Kid Gall. Eso. Liver Stom. Aorta IVC Panc. RAG LAG Duo. Blad. Pros. Avg nn-UNet (350 Ep) 0.967 0.855 0.958 0.663 0.736 0.973 0.888 0.956 0.907 0.793 0.533 0.572 0.668 - - 0.805 nn-UNet (1000 Ep) 0.973 0.940 0.965 0.681 0.810 0.980 0.893 0.967 0.917 0.834 0.667 0.689 0.701 - - 0.847 TransBTS 0.956 0.957 0.955 0.619 0.770 0.974 0.867 0.958 0.852 0.836 0.591 0.630 0.648 - - 0.816 UNETR 0.942 0.956 0.930 0.552 0.741 0.967 0.836 0.947 0.829 0.815 0.564 0.621 0.624 - - 0.794 nnFormer 0.949 0.952 0.950 0.601 0.758 0.972 0.859 0.960 0.843 0.832 0.569 0.618 0.637 - - 0.808 SwinUNETR 0.972 0.961 0.961 0.649 0.814 0.978 0.889 0.961 0.862 0.854 0.659 0.649 0.664 - - 0.836 3D UX-Net (k=7) 0.971 0.965 0.966 0.603 0.828 0.978 0.869 0.962 0.878 0.837 0.696 0.689 0.696 - - 0.841 3D UX-Net (k=21) 0.968 0.962 0.967 0.610 0.830 0.977 0.858 0.954 0.880 0.829 0.701 0.697 0.700 - - 0.840 UNesT-B 0.971 0.965 0.967 0.615 0.831 0.980 0.865 0.949 0.883 0.845 0.691 0.700 0.697 - - 0.843 RepOptimizer 0.970 0.967 0.971 0.635 0.823 0.978 0.875 0.963 0.882 0.850 0.689 0.691 0.711 - - 0.847 Rep3D (Fixed) 0.972 0.965 0.970 0.644 0.838 0.980 0.883 0.965 0.893 0.861 0.714 0.701 0.725 - - 0.855 Rep3D (LRBM) 0.975 0.969 0.975 0.657 0.845 0.984 0.891 0.970 0.901 0.879 0.718 0.721 0.750 - - 0.864*
@@ 5.1 Evaluation on Tissue & Tumor Segmentation
To assess the generalization and scalability of Rep3D across diverse anatomical structures and clinical targets, we evaluate performance on three representative volumetric segmentation tasks using the KiTS, MSD Pancreas, and MSD Hepatic Vessel datasets. As shown in Table 1, Rep3D achieves state-of-the-art performance across all settings, consistently outperforming both convolution- and transformer-based baselines. On the KiTS dataset, which includes kidney, tumor, and cyst segmentation, Rep3D achieves the highest average Dice score of 0.736, with strong individual scores of 0.955 (kidney), 0.763 (tumor), and 0.490 (cyst). Notably, Rep3D improves tumor segmentation performance by 2.28% Dice over UNesT-B and 5.39% Dice over 3D UX-Net, demonstrating its ability to adapt to complex local variations in pathological regions. On the MSD Pancreas task, which is particularly challenging due to the pancreas’s low contrast and irregular boundaries, Rep3D sets a new benchmark with an average Dice score of 0.723, outperforming SwinUNETR (0.708), nnUNet (0.703), and UNesT-B (0.690). Tumor segmentation also benefits from our re-parameterization design, improving by 3.32% Dice compared to 3D UX-Net and 2.03% Dice compared to the fixed-prior variant. On the MSD Hepatic Vessel dataset, Rep3D continues to lead with a mean Dice of 0.674, outperforming the previous best model (UNesT-B, 0.640) and demonstrating superior vessel and tumor localization. The results also highlight the effectiveness of Rep3D’s spatially adaptive learning dynamics, especially in sparse and small-structure segmentation where traditional large-kernel convolutions or global self-attention tend to underperform.
Table 3: Ablation Studies on bias generator’s convolutional layers and LRBM 3D adaptability.
[…]